OpenStudy (anonymous):

Proofs: prove if |s|<=t ^ lim(t)=0, then lim(s)=0

terenzreignz (terenzreignz):

s and t are functions?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

Well, squeeze.

terenzreignz (terenzreignz):

If \(\large |s|\le t\) then either... \[\large -t \le s \le t\] or \[\large t \le s \le -t\] Either way, squeeze.

terenzreignz (terenzreignz):

ON second thought, the "second case" \[\Large \color{red}{t\le s \le -t}\] is not possible So just stick to this statement... \[\huge \color{blue}{-t \le s \le t}\] and use squeeze